OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

help please is the answer x=2 correct me if I'm wrong please ?

OpenStudy (anonymous):

2x + y = 1 (Eqn 1) 4x + 2y = -2 Dividing by 2: 2x+y=-1 (Eqn 2) 2x + y = 1 2x + y= -1 Dividing here will yield no solution.

OpenStudy (anonymous):

so is the answer c ?

OpenStudy (anonymous):

thank you btw

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks so much